# DG Singular equivalence and singular locus

### v2

@@ 3.2. Hochschild cohomology of DG singularity category
$\mathbb{R}\mathrm{Hom}_{R}(R,J^{*})\rightarrow\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),J^{*})$
induced by Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$. Since $\mathbb{R}\mathrm{Hom}_{R}(R,J^{*})$ is a bounded finitely generated $R^{e}$-module complex, it admits a finitely generated projective $R^{e}$-module resolution, denoted by $P^{\ast}_{J^{*}}$. It suggests that there are quasi-isomorphisms
$\mathrm{Bar}(R)\otimes_{R^{e}}P^{\ast}_{J^{*}}\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R}(R,J^{*})\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),J^{*})$
$\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R}(\mathrm{Bar}(R),R^{e})\xrightarrow{\sim}E^{*}.$
@@ Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions. Moreover, we appreciate Martin Kalck for pointing out an error in the previous version of this paper.

### v5

@@ 3.2. Hochschild cohomology of DG singularity category
$\mathbb{R}\mathrm{Hom}_{R^{e}}(R,J^{*})\rightarrow\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$
induced by Bar resolution $\mathrm{Bar}(R)\twoheadrightarrow R$. Due to the boundedness of $J^{\ast}$, there are only finite many nontrivial cohomologies of $\mathbb{R}\mathrm{Hom}_{R^{e}}(R,J^{*})$.
In the meantime, for any $i\in\mathbb{Z}$, $\mathrm{Hom}_{K(R^{e})}(P^{*}_{R},R^{e}[i])\cong\mathrm{Hom}_{D(R^{e})}(R,R^{e}[i])$ as $R^{e}$-modules holds for any finitely generated projective $R^{e}$-module resolution $P^{*}_{R}$ of $R$, where $K(R^{e})$ is the homotopy category of complexes of $R^{e}$-modules. Hence we have that for any $i\in\mathbb{Z}$, $\mathrm{Hom}_{D(R^{e})}(R,R^{e}[i])$ is a finitely generated $R^{e}$-module.
By [12, Proposition 3.5], we get that $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$ is quasi-isomorphic to a bounded finitely generated $R^{e}$-module complex. Thus, $\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$ admits a finitely generated projective $R^{e}$-module resolution, denoted by $P^{\ast}_{J^{*}}$. It suggests that there are quasi-isomorphisms
$\mathrm{Bar}(R)\otimes_{R^{e}}P^{\ast}_{J^{*}}\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}(R,J^{*})\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),J^{*})$
$\xrightarrow{\sim}\mathrm{Bar}(R)\otimes_{R^{e}}\mathbb{R}\mathrm{Hom}_{R^{e}}(\mathrm{Bar}(R),R^{e})\xrightarrow{\sim}E^{*}.$
@@ Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions. Moreover, we thank Martin Kalck for pointing out an error in the previous version of this paper.